OpenStudy (anonymous):

what is the answer to pi/2 minus pi/7

OpenStudy (anonymous):

same as \[\frac{1}{2}-\frac{1}{7}\] just stick a \(\pi\) next to it

OpenStudy (anonymous):

create a common denominator \[\frac{\pi}{2} *\frac{7}{7}-\frac{\pi}{7} *\frac{2}{2}\]

OpenStudy (anonymous):

7pi/14 - 2pi/14 gives you 5pi/14